OpenStudy (anonymous):

Please help me with these two problems :( a + b = 11 a – b = 27 20c - 30d = 70 20c - 8d = -32

OpenStudy (hoblos):

a + b = 11 a – b = 27 add the two equation to get 2a=38 --> a=19 b=11-a=11-19=-8

OpenStudy (hoblos):

can you do the second one by yourself now?

OpenStudy (anonymous):

WELL if I could have I would have done it myself.

OpenStudy (mysesshou):

I'll help walk you through it if you want. :)

OpenStudy (mysesshou):

20c - 30d = 70 20c - 8d = -32 First, I would multiply through by -1 to one of the equations. So what would that give you ?

OpenStudy (mysesshou):

20c - 30d = 70 (eq1) -1 * (20c - 8d) = -1*(-32) (eq2) By multiplying eq2 by -1, then you are able to add the equations together and the C is eliminated. ( i suppose you could have just subtracted them, but adding is sometimes simpler) -20c + 8d = 32 (eq2)

OpenStudy (mysesshou):

Can you take the new eq2 and add it to the original eq1 ? Add these. 20c - 20c = 0 20c - 30d = 70 (eq1) -20c + 8d = 32 (eq2) What would that give you?